OpenStudy (anonymous):

Company A rents copiers for a monthly charge of $200 plus 10 cents per copy. Company B rents copiers for a monthly charge of $400 plus 5 cents per copy. What is the number of copies above which Company A's charges are the higher of the two?

OpenStudy (anonymous):

Company A=200+ 0.1x Company B=400+0.05x 200+ 0.1(x-400)=400+0.05(x-200) 200+0.1x-40=400+0.05x-10 160+ 0.1x = 390 +0.05x 0.1x+0.05x = 390-160 0.05x = 230 x = $4.600